3.10 The following data give the 2014 profits (in millions of dollars) of the top 10 companies listed in the 2014 Fortune 500 (source: www.fortune.com). 2014 Profits Company Wal-Mart Stores Exxon Mobil Chevron Berkshire Hathaway Apple Phillips 66 General Motors Ford Motor General Electric Valero Energy (millions of dollars) 16,022 32,580 21,423 19,476 37,037 3726 5346 7155 13,057 2720 Find the mean and median for these data. Do these data have a mode? Explain
Answer #1

putting above data in ascending order:

s. no x
1 2720
2 3726
3 5346
4 7155
5 13057
6 16022
7 19476
8 21423
9 32580
10 37037

mean =total/n= 158542/10=15854.2

median =middle value =(13057+16022)/2=14539.5

there is no mode as all values appears only once; therefore there is no value is most frequent,